You (safely) bungee jump from a 200-feet tall bridge in your town

1. You (safely) bungee jump from a 200-feet tall bridge in your town. Your distance above the water's surface depends on the time since you jumped. Sketch a reasonable graph.
Answer #1

$$ y=y_{i}+u t+\frac{1}{2} a t^{2} $$ 

\(y_{i}=\) Pnitial positim \(=800\) feet 

\(u=i\) utial speed \(=0\) 

\(a=a\) cceleration \(=-32 \cdot 2 \mathrm{ft} / \mathrm{sec}^{2}\) 

$$ \begin{array}{l} t= time \\ y=200-\frac{1}{2} \times 32 \cdot 2 t^{2} \end{array} $$ 

\(\Rightarrow y=200-16 \cdot 1 t^{2}\) t can be replace with x 

so, \( y=200-16 \cdot 1 x^{2}.\) 

So, the position as a function of time will be parabola starting at 200 pointing downward and hitting the \(x\) -axis at \(y=0\)

image.png

answered by: Pixies